Complete the square of the terms inside the square brackets:


ax2 + bx + c = a [x2 + ( b/a) x + c/a]




1
Expert's answer
2021-05-07T11:42:25-0400

"ax^2+bx+c=a\Big[x^2+\Big(\dfrac{b}{a}\Big)x+\dfrac{c}{a}\Big]"

"=a\Big[x^2+\Big(\dfrac{b}{a}\Big)x+\dfrac{c}{a}+\dfrac{b^2}{4a^2}-\dfrac{b^2}{4a^2}\Big]"

"=a\Big[\Big(x^2+\Big(\dfrac{b}{a}\Big)x+\dfrac{b^2}{4a^2}\Big)-\dfrac{b^2}{4a^2}+\dfrac{c}{a}\Big]"

"=a\Big[\Big(x+\dfrac{b}{2a}\Big)^2-\dfrac{b^2}{4a^2}+\dfrac{c}{a}\Big]"

"=a\Big[x+\dfrac{b}{2a}\Big]^2-\dfrac{b^2}{4a}+c"